# Large Language Models for Information Retrieval: A Survey

### v2

@@ 4.2 Employing LLMs to Enhance Model Architecture
Leveraging the excellent text encoding and decoding capabilities of LLMs, it is feasible to understand queries and documents with greater precision compared to earlier smaller-sized models [59]. Researchers have endeavored to utilize LLMs as the foundation for constructing advanced retrieval models. These methods can be grouped into two categories, i.e., encoder-based retrievers and generative retrievers.
4.2.1 Encoder-based Retriever
@@
General Retriever. Since the effectiveness of retrievers primarily relies on the capability of text embedding, the evolution of text embedding models often has a significant impact on the progress of retriever development. In the era of LLMs, a pioneer work is made by OpenAI [115]. They view the adjacent text segments as positive pairs to facilitate the unsupervised pre-training of a set of text embedding models, denoted as cpt-text, whose parameter values vary from 300M to 175B. Experiments conducted on the MS MARCO [102] and BEIR [114] datasets indicate that larger model scales have the potential to yield improved performance in unsupervised learning and transfer learning for text search tasks. Nevertheless, pre-training LLMs from scratch is prohibitively expensive for most researchers. To overcome this limitation, T5-family models with smaller sizes, such as Base, Large, XL, and XXL, are used to initialize the model parameters of bi-encoders, which are then fine-tuned on retrieval datasets [116]. The experimental results confirm again that larger model sizes can lead to better performance, particularly in zero-shot settings.
Task-aware Retriever. While the aforementioned studies primarily focus on using LLMs as text embedding models for downstream retrieval tasks, retrieval performance can be greatly enhanced when task-specific instructions are integrated. For example, TART [117] devises a task-aware retrieval model that introduces a task-specific instruction before the question. This instruction includes descriptions of the task’s intent, domain, and desired retrieved unit. For instance, given that the task is question-answering, an effective prompt might be “Retrieve a Wikipedia text that answers this question. {question}”. Here, “Wikipedia” (domain) indicates the expected source of retrieved documents, “text” (unit) suggests the type of content to retrieve, and “answers this question” (intent) demonstrates the intended relationship between the retrieved texts and the question. This approach can take advantage of the powerful language modeling capability and extensive knowledge of LLMs to precisely capture the user’s search intents across various retrieval tasks.
@@ 4.2.2 Generative Retriever
Traditional IR systems typically follow the “index-retrieval-rank” paradigm to locate relevant documents based on user queries, which has proven effective in practice. However, these systems usually consist of three separate modules: the index module, the retrieval module, and the reranking module. Therefore, optimizing these modules collectively can be challenging, potentially resulting in sub-optimal retrieval outcomes. Additionally, this paradigm demands additional space for storing pre-built indexes, further burdening storage resources. Recently, model-based generative retrieval methods [118, 119, 120] have emerged to address these challenges. These methods move away from the traditional “index-retrieval-rank” paradigm and instead use a unified model to directly generate document identifiers (i.e., DocIDs) relevant to the queries. In these model-based generative retrieval methods, the knowledge of document corpus is stored in the model parameters, eliminating the need for additional storage space for the index. Existing methods have explored generating document identifiers through fine-tuning and prompting of LLMs [121, 122]
@@
TABLE IV: The comparison of retrievers that leverage LLMs as the foundation.
@@
cpt-text [115] cpt-text Encoder-based Training from scratch
GTR [116] T5 Encoder-based Fine-tuning
TART [117] T5 Encoder-based
@@ 5 Reranker
Reranker, as the second-pass document filter in IR, aims to rerank a document list retrieved by the retriever (e.g., BM25) based on the query-document relevance. With respect to the application of LLMs, the existing reranking methods can be divided into three primary paradigms: fine-tuning LLMs for reranking, prompting LLMs for reranking, and utilizing LLMs for training data augmentation. These paradigms will be elaborated upon in the following sections. Recall that we will use the term document to refer to the text retrieved in general IR scenarios, including instances such as passages (e.g., passages in MS MARCO passage ranking dataset [102]).
5.1 Fine-tuning LLMs for Reranking
@@
Fine-tuning is an important step in applying pre-trained LLMs to a reranking task. Due to the lack of awareness of ranking during pre-training, LLMs cannot appropriately measure the query-document relevance and fully understand the reranking tasks. By fine-tuning LLMs on task-specific ranking datasets, such as the MS MARCO passage ranking dataset [102], which includes signals of both relevance and irrelevance, LLMs can adjust their parameters to yield better performance in the reranking tasks. In general, existing fine-tuning strategies for LLMs can be categorized into two principal approaches: (1) fine-tuning LLMs as generation models and (2) fine-tuning LLMs as ranking models.
5.1.1 Fine-tuning LLMs as Generation Models
@@
In this field, existing studies usually formulate document ranking as a generation task and optimize LLMs by generative loss [13, 123, 124]. Specifically, reranking models are usually fine-tuned to generate a single token, such as “true” or “false”, given the query and documents. During inference, the query-document relevance score is determined based on the logit of the generated token. For example, a T5 model can be fine-tuned to generate classification tokens for relevant or irrelevant query-document pairs [13]. At inference time, a softmax function is applied to the logits of “true” and “false” tokens, and the relevance score is calculated as the probability of the “true” token. The following method [123] involves a multi-view learning approach based on the T5 model. This approach simultaneously considers two tasks: generating classification tokens for a given query-document pair and generating the corresponding query conditioned on the provided document. DuoT5 [124] considers a triple $(q,d_{i},d_{j})$ as the input of the T5 model and is fine-tuned to generate token “true” if document $d_{i}$ is more relevant to query $q_{i}$ than document $d_{j}$, and “false” otherwise. During inference, for each document $d_{i}$, it enumerates all other documents $d_{j}$ and uses global aggregation functions to generate the relevance score $s_{i}$ for document $d_{i}$ (e.g., $s_{i}=\sum_{j}p_{i,j}$, where $p_{i,j}$ represents the probability of generating “true” when taking $(q,d_{i},d_{j})$ as the model input).
5.1.2 Fine-tuning LLMs as Ranking Models
@@
Although the methods of fine-tuning LLMs as generation models outperform some strong ranking baselines, they are not optimal for reranking tasks. This stems from two primary reasons. First, it is commonly expected that a reranking model will yield a numerical relevance score for each query-document pair rather than text tokens. Second, compared to generation losses, it is more reasonable to optimize the reranking model using ranking losses (e.g., RankNet [125]). While some pre-trained models, such as BERT, have been explored for document reranking, the use of seq2seq-based LLMs, such as T5-3B, for reranking tasks has not yet been thoroughly investigated. Recently, RankT5 [126] directly calculates the relevance score for a query-document pair and optimizing the ranking performance with “pairwise” or “listwise” ranking losses. An avenue for potential performance enhancement lies in the substitution of the base-sized T5 model with its larger-scale counterpart.
@@
Figure 6: Three types of prompting-based reranking methods: (a) pointwise methods that consist of relevance generation (upper) and query generation (lower), (b) listwise methods, and (c) pairwise methods.
5.2 Prompting LLMs for Reranking
@@
As the size of LLMs scales up (e.g., exceeding 10 billion parameters), it becomes increasingly difficult to fine-tune the reranking model. Addressing this challenge, recent efforts have attempted to leverage the exceptional instructional capabilities of LLMs to directly enhance document reranking via prompting strategies. In general, these prompting strategies for reranking can be divided into three categories: pointwise, listwise, and pairwise methods. A comprehensive exploration of these strategies follows in the subsequent sections.
@@ 5.2.1 Pointwise methods
The pointwise methods measure the relevance between a query and a single document, and can be categorized into two types: relevance generation [127] and query generation [128, 129]. The upper part in Figure 6 (a) shows the relevance generation based on a given prompt, where LLMs output “Yes” if the document is relevant to the query and “No” otherwise. The probability that LLMs generate the token “Yes” or “No” is used to calculate the query-document relevance score $s$ as follows:
$\text{score}=\begin{cases}1+p(\text{Yes}),&\text{if LLMs output Yes}\\ 1-p(\text{No}),&\text{if LLMs output No}\end{cases}$ (1)
where $p(\text{Yes/No})$ represents the generation probability of “Yes” or “No” token.
As for the query generation, the lower part in Figure 6 (a) shows an example in which the query-document relevance score is determined by the average log-likelihood of generating the actual query tokens based on the document:
@@
where $|q|$ denotes the token number of query $q$, $d$ denotes the document, and $\mathcal{P}$ represents the provided prompt. The documents are then reranked based on their relevance scores. Existing query generation-based methods [128] primarily rely on a handcrafted prompt (e.g., “Please write a query based on this document”), which may not be optimal. As prompt is a key factor in instructing LLMs to perform various NLP tasks, it is important to identify the optimal prompt for best model performance. Along this line, a discrete prompt optimization method Co-Prompt [129] is proposed for better prompt generation in reranking tasks.
@@ 5.2.2 Listwise Methods
Pointwise methods calculate the query-document relevance score using the log probability of output tokens, which is usually unavailable for LLM APIs (e.g., ChatGPT and GPT-4). Recently, listwise methods [130, 131] that directly rank a list of documents have been proposed as a solution to this problem. These methods insert the query and a document list into the prompt and instruct the LLMs to output the reranked document identifiers (see Figure 6 (b)). Considering the limited input length of LLMs, these methods also employ a sliding window strategy to rerank a subset of candidate documents each time [130]. Extensive experiments based on several LLMs reveal that the GPT-4-based method achieves competitive results and even outperforms some supervised methods on some IR benchmarks.
@@ 5.2.3 Pairwise Methods
Although listwise methods have yielded promising performance, they still suffer from some weaknesses. First, according to the experimental results [130], only the GPT-4-based method is able to achieve competitive performance. When using smaller parameterized language models (e.g., FLAN-UL2 with 20B parameters), listwise methods may produce very few usable results and underperform many supervised methods. Second, the performance of listwise methods is highly sensitive to the document order in the prompt. When the document order is randomly shuffled, listwise methods perform even worse than BM25. These weaknesses may be due to the fact that existing popular LLMs are generally not pre-trained against ranking tasks.
A recent study [132] has shown that LLMs inherently have a sense of pairwise document comparison, which is much simpler than directly outputting a reranked document list. In this study, a pairwise ranking prompt (see Figure 6 (c)) is first designed to compare whether one document is more relevant than another document. Then, several ranking algorithms are devised to leverage this pairwise ranking prompt as the computation unit for reranking the whole document list. The experimental results show the state-of-the-art performance on the standard benchmarks using moderate-size LLMs (e.g., Flan-UL2 with 20B parameters), which are much smaller than GPT-4.

### v3

@@ 4.2 Employing LLMs to Enhance Model Architecture
Leveraging the excellent text encoding and decoding capabilities of LLMs, it is feasible to understand queries and documents with greater precision compared to earlier smaller-sized models [59]. Researchers have endeavored to utilize LLMs as the foundation for constructing advanced retrieval models. These methods can be grouped into two categories, i.e., dense retrievers and generative retrievers.
4.2.1 Dense Retriever
@@
General Retriever. Since the effectiveness of retrievers primarily relies on the capability of text embedding, the evolution of text embedding models often has a significant impact on the progress of retriever development. In the era of LLMs, a pioneer work is made by OpenAI [130]. They view the adjacent text segments as positive pairs to facilitate the unsupervised pre-training of a set of text embedding models, denoted as cpt-text, whose parameter values vary from 300M to 175B. Experiments conducted on the MS MARCO [111] and BEIR [128] datasets indicate that larger model scales have the potential to yield improved performance in unsupervised learning and transfer learning for text search tasks. Nevertheless, pre-training LLMs from scratch is prohibitively expensive for most researchers. To overcome this limitation, some studies [131, 133] use pre-trained LLMs to initialize the bi-encoder of dense retriever. Specifically, GTR [133] adopts T5-family models, including T5-base, Large, XL, and XXL, to initialize and fine-tune dense retrievers. RepLLaMA [131] further fine-tunes the LLaMA model on multiple stages of IR, including retrieval and reranking. For the dense retrieval task, RepLLaMA appends an end-of-sequence token “<</s>>” to the input sequences, i.e., queries or documents, and regards its output embeddings as the representation of queries or documents. The experiments confirm again that larger model sizes can lead to better performance, particularly in zero-shot settings. Notably, the researchers of RepLLaMA [131] also study the effectiveness of applying LLaMA in the reranking stage, which will be introduced in Section 5.1.3.
Task-aware Retriever. While the aforementioned studies primarily focus on using LLMs as text embedding models for downstream retrieval tasks, retrieval performance can be greatly enhanced when task-specific instructions are integrated. For example, TART [132] devises a task-aware retrieval model that introduces a task-specific instruction before the question. This instruction includes descriptions of the task’s intent, domain, and desired retrieved unit. For instance, given that the task is question-answering, an effective prompt might be “Retrieve a Wikipedia text that answers this question. {question}”. Here, “Wikipedia” (domain) indicates the expected source of retrieved documents, “text” (unit) suggests the type of content to retrieve, and “answers this question” (intent) demonstrates the intended relationship between the retrieved texts and the question. This approach can take advantage of the powerful language modeling capability and extensive knowledge of LLMs to precisely capture the user’s search intents across various retrieval tasks. Considering the efficiency of retrievers, it first fine-tunes a TART-full model with cross-encoder architecture, which is initialized from LLMs (e.g., T0-3B, Flan-T5). Then, a TART-dull model initialized from Contriever [134] is learned by distillating knowledge from the TART-full.
@@ 4.2.2 Generative Retriever
Traditional IR systems typically follow the “index-retrieval-rank” paradigm to locate relevant documents based on user queries, which has proven effective in practice. However, these systems usually consist of three separate modules: the index module, the retrieval module, and the reranking module. Therefore, optimizing these modules collectively can be challenging, potentially resulting in sub-optimal retrieval outcomes. Additionally, this paradigm demands additional space for storing pre-built indexes, further burdening storage resources. Recently, model-based generative retrieval methods [135, 136, 137] have emerged to address these challenges. These methods move away from the traditional “index-retrieval-rank” paradigm and instead use a unified model to directly generate document identifiers (i.e., DocIDs) relevant to the queries. In these model-based generative retrieval methods, the knowledge of the document corpus is stored in the model parameters, eliminating the need for additional storage space for the index. Existing methods have explored generating document identifiers through fine-tuning and prompting of LLMs [138, 139]
@@
TABLE IV: The comparison of retrievers that leverage LLMs as the foundation. “KD” is short for “Knowledge Distillation”.
@@
cpt-text [130] GPT-series Dense
Pre-training
Fine-tuning
GTR [133] T5 Dense
Pre-training &
Fine-tuning
RepLLaMA [131] LLAMA Dense Fine-tuning
TART-full [132]
T0 &
Flan-T5
Dense
@@ 5 Reranker
Reranker, as the second-pass document filter in IR, aims to rerank a document list retrieved by the retriever (e.g., BM25) based on the query-document relevance. Based on the usage of LLMs, the existing LLM-based reranking methods can be divided into three paradigms: utilizing LLMs as supervised rerankers, utilizing LLMs as unsupervised rerankers, and utilizing LLMs for training data augmentation. These paradigms are summarized in Table V and will be elaborated upon in the following sections. Recall that we will use the term document to refer to the text retrieved in general IR scenarios, including instances such as passages (e.g., passages in MS MARCO passage ranking dataset [111]).
TABLE V: Summary of existing LLM-based re-ranking methods. “Enc” and “Dec” denote encoder and decoder, respectively.
Paradigm Type Method
Supervised
Rerankers
Enc-only [140]
Enc-dec [13], [141], [142], [143]
Dec-only [131], [144], [145]
Unsupervised
Rerankers
Pointwise [146], [147], [148], [149], [150], [151]
Listwise [152], [153], [154]
Pairwise [155], [156]
Data Augmentation - [157], [158], [159], [160], [161], [162]
5.1 Utilizing LLMs as Supervised Rerankers
@@
Supervised fine-tuning is an important step in applying pre-trained LLMs to a reranking task. Due to the lack of awareness of ranking during pre-training, LLMs cannot appropriately measure the query-document relevance and fully understand the reranking tasks. By fine-tuning LLMs on task-specific ranking datasets, such as the MS MARCO passage ranking dataset [111], which includes signals of both relevance and irrelevance, LLMs can adjust their parameters to yield better performance in the reranking tasks. Based on the backbone model structure, we can categorize existing supervised rerankers as: (1) encoder-only, (2) encoder-decoder, and (3) decoder-only.
5.1.1 Encoder-only
The encoder-based rerankers represent a significant turning point in applying LLMs to document ranking tasks. They demonstrate how some pre-trained language models (e.g., BERT [59]) can be finetuned to deliver highly accurate relevance predictions. A representative approach is monoBERT [140], which transforms a query-document pair into a sequence “[CLS] query [SEP] document [SEP]” as the model input and calculates the relevance score by feeding the “[CLS]” representation into a linear layer. The reranking model is optimized based on the cross-entropy loss.
5.1.2 Encoder-Decoder
@@
In this field, existing studies mainly formulate document ranking as a generation task and optimize an encoder-decoder-based reranking model [13, 141, 142, 143]. Specifically, given the query and the document, reranking models are usually fine-tuned to generate a single token, such as “true” or “false”. During inference, the query-document relevance score is determined based on the logit of the generated token. For example, a T5 model can be fine-tuned to generate classification tokens for relevant or irrelevant query-document pairs [13]. At inference time, a softmax function is applied to the logits of “true” and “false” tokens, and the relevance score is calculated as the probability of the “true” token. The following method [141] involves a multi-view learning approach based on the T5 model. This approach simultaneously considers two tasks: generating classification tokens for a given query-document pair and generating the corresponding query conditioned on the provided document. DuoT5 [142] considers a triple $(q,d_{i},d_{j})$ as the input of the T5 model and is fine-tuned to generate token “true” if document $d_{i}$ is more relevant to query $q_{i}$ than document $d_{j}$, and “false” otherwise. During inference, for each document $d_{i}$, it enumerates all other documents $d_{j}$ and uses global aggregation functions to generate the relevance score $s_{i}$ for document $d_{i}$ (e.g., $s_{i}=\sum_{j}p_{i,j}$, where $p_{i,j}$ represents the probability of generating “true” when taking $(q,d_{i},d_{j})$ as the model input).
@@
Although these generative loss-based methods outperform several strong ranking baselines, they are not optimal for reranking tasks. This stems from two primary reasons. First, it is commonly expected that a reranking model will yield a numerical relevance score for each query-document pair rather than text tokens. Second, compared to generation losses, it is more reasonable to optimize the reranking model using ranking losses (e.g., RankNet [163]). Recently, RankT5 [143] has directly calculated the relevance score for a query-document pair and optimized the ranking performance with “pairwise” or “listwise” ranking losses. An avenue for potential performance enhancement lies in the substitution of the base-sized T5 model with its larger-scale counterpart.
5.1.3 Decoder-only
Recently, there have been some attempts [131, 144, 145] to rerank documents by fine-tuning decoder-only models (such as LLaMA). For example, RankLLaMA [131] proposes formatting the query-document pair into a prompt “query: {query} document: {document} [EOS]” and utilizes the last token representation for relevance calculation. Besides, RankingGPT [144] has been proposed to bridge the gap between LLMs’ conventional training objectives and the specific needs of document ranking through two-stage training. The first stage involves continuously pretraining LLMs using a large number of relevant text pairs collected from web resources, helping the LLMs to naturally generate queries relevant to the input document. The second stage focuses on improving the model’s text ranking performance using high-quality supervised data and well-designed loss functions. Different from these pointwise rerankers [131, 144], Rank-without-GPT [145] proposes to train a listwise reranker that directly outputs a reranked document list. The authors first demonstrate that existing pointwise datasets (such as MS MARCO [111]), which only contain binary query-document labels, are insufficient for training efficient listwise rerankers. Then, they propose to use the ranking results of existing ranking systems (such as Cohere rerank API) as gold rankings to train a listwise reranker based on Code-LLaMA-Instruct.
@@ 5.1.3 Decoder-only
Figure 6: Three types of unsupervised reranking methods: (a) pointwise methods that consist of relevance generation (upper) and query generation (lower), (b) listwise methods, and (c) pairwise methods.
5.2 Utilizing LLMs as Unsupervised Rerankers
@@
As the size of LLMs scales up (e.g., exceeding 10 billion parameters), it becomes increasingly difficult to fine-tune the reranking model. Addressing this challenge, recent efforts have attempted to prompt LLMs to directly enhance document reranking in an unsupervised way. In general, these prompting strategies can be divided into three categories: pointwise, listwise, and pairwise methods. A comprehensive exploration of these strategies follows in the subsequent sections.
@@ 5.2.1 Pointwise methods
The pointwise methods measure the relevance between a query and a single document, and can be categorized into two types: relevance generation [146, 147] and query generation [148, 149, 150].
The upper part in Figure 6 (a) shows an example of relevance generation based on a given prompt, where LLMs output a binary label (“Yes” or “No”) based on whether the document is relevant to the query. Following [13], the query-document relevance score $f(q,d)$ can be calculated based on the log-likelihood of token “Yes” and “No” with a softmax function:
$f(q,d)=\frac{\text{exp}(S_{Y})}{\text{exp}(S_{Y})+\text{exp}(S_{N})},$ (1)
where $S_{Y}$ and $S_{N}$ represent the LLM’s log-likelihood scores of “Yes” and “No” respectively. In addition to binary labels, Zhuang et al. [147] propose to incorporate fine-grained relevance labels (e.g., “highly relevant”, “somewhat relevant” and “not relevant”) into the prompt, which helps LLMs more effectively differentiate among documents with varying levels of relevance to a query.
As for the query generation shown in the lower part of Figure 6 (a), the query-document relevance score is determined by the average log-likelihood of generating the actual query tokens based on the document:
@@
where $|q|$ denotes the token number of query $q$, $d$ denotes the document, and $\mathcal{P}$ represents the provided prompt. The documents are then reranked based on their relevance scores. It has been proven that some LLMs (such as T0) yield significant performance in zero-shot document reranking based on the query generation method [148]. Recently, research [149] has also shown that the LLMs that are pre-trained without any supervised instruction fine-tuning (such as LLaMA) also yield robust zero-shot ranking ability.
Although effective, these methods primarily rely on a handcrafted prompt (e.g., “Please write a query based on this document”), which may not be optimal. As prompt is a key factor in instructing LLMs to perform various NLP tasks, it is important to optimize prompt for better performance. Along this line, a discrete prompt optimization method Co-Prompt [150] is proposed for better prompt generation in reranking tasks. Besides, PaRaDe [151] proposes a difficulty-based method to select few-show demonstrations to include in the prompt, proving significant improvements compared with zero-shot prompts.
Note that these pointwise methods rely on accessing the output logits of LLMs to calculate the query-document relevance scores. As a result, they are not applicable to closed-sourced LLMs, whose API-returned results do not include logits.
TABLE VI: The comparison between different methods. NN denotes the number of documents to rerank. The Complexity, Logits, and Batch represent the computational complexity, whether accesses LLM’s logits, and whether allows batch inference respectively. $k$ is the constant in sliding windows strategy. As for the Performance, we use NDCG@10 as a metric, and the results are calculated by reranking the top 100 documents retrieved by BM25 on TREC-DL2019 and TREC-DL2020. The best model is in bold while the second-best is marked with an underline. The results come from previous study [155]. *Since the parameters of ChatGPT have not been released, its model parameters are based on public estimates [164].
Methods LLM Size Properties Performance
Complexity Logits Batching TREC-DL19 TREC-DL20
Initial Retriever BM25 - - - - - 50.58 47.96
Supervised monoBERT [140] BERT 340M - $\checkmark$ $\checkmark$ 70.50 67.28
monoT5 [13] T5 220M - $\checkmark$ $\checkmark$ 71.48 66.99
RankT5 [143] T5 3B - $\checkmark$ $\checkmark$ 71.22 69.49
Unsupervised-Pointwise Query Generation [148] FLAN-UL2 20B $O(N)$ $\checkmark$ $\checkmark$ 58.95 60.02
Relevance Generation [146] FLAN-UL2 20B $O(N)$ $\checkmark$ $\checkmark$ 64.61 65.39
Unsupervised-Listwise RankGPT3.5 [152] gpt-3.5-turbo 154B* O(kN)O(k*N) 65.80 62.91
RankGPT4 [152] gpt-4 1T* O(kN)O(k*N) 75.59 70.56
Unsupervised-Pairwise PRP-Allpair [155] FLAN-UL2 20B $O(N^{2})$ $\checkmark$ $\checkmark$ 72.42 70.68
PRP-Heapsort [155] FLAN-UL2 20B O(NlogN)O(N*logN) $\checkmark$ 71.88 69.43
@@ 5.2.2 Listwise Methods
Listwise methods [152, 153] aim to directly rank a list of documents (see Figure 6 (b)). These methods insert the query and a document list into the prompt and instruct the LLMs to output the reranked document identifiers. Due to the limited input length of LLMs, it is not feasible to insert all candidate documents into the prompt. To alleviate this issue, these methods employ a sliding window strategy to rerank a subset of candidate documents each time. This strategy involves ranking from back to front using a sliding window, re-ranking only the documents within the window at a time.
Although listwise methods have yielded promising performance, they still suffer from some weaknesses. First, according to the experimental results [152], only the GPT-4-based method can achieve competitive performance. When using smaller parameterized language models (e.g., FLAN-UL2 with 20B parameters), listwise methods may produce very few usable results and underperform many supervised methods. Second, the performance of listwise methods is highly sensitive to the document order in the prompt. When the document order is randomly shuffled, listwise methods perform even worse than BM25 [152], revealing positional bias issues in the listwise ranking of LLMs. To alleviate this issue, Tang et al. [154] introduce a permutation self-consistency method, which involves shuffling the list in the prompt and aggregating the generated results to achieve a more accurate and unbiased ranking.
@@ 5.2.3 Pairwise Methods
In pairwise methods [155], LLMs are given a prompt that consists of a query and a document pair (see Figure 6 (c)). Then, they are instructed to generate the identifier of the document with higher relevance. To rerank all candidate documents, aggregation methods like AllPairs are used. AllPairs first generates all possible document pairs and aggregates a final relevance score for each document. To speed up the ranking process, efficient sorting algorithms, such as heap sort and bubble sort, are usually employed [155]. These sorting algorithms utilize efficient data structures to compare document pairs selectively and elevate the most relevant documents to the top of the ranking list, which is particularly useful in top-$k$ ranking. Experimental results show the state-of-the-art performance on the standard benchmarks using moderate-size LLMs (e.g., Flan-UL2 with 20B parameters), which are much smaller than those typically employed in listwise methods (e.g., GPT3.5).
Although effective, pairwise methods still suffer from high time complexity. To alleviate the efficiency problem, a setwise approach [156] has been proposed to compare a set of documents at a time and select the most relevant one from them. This approach allows the sorting algorithms (such as heap sort) to compare more than two documents at each step, thereby reducing the total number of comparisons and speeding up the sorting process.
5.2.4 Comparison and Discussion
In this part, we will compare different unsupervised methods from various aspects to better illustrate the strengths and weaknesses of each method, which is summarized in Table VI. We choose representative methods [148, 146, 152, 155] in pointwise, listwise and pairwise ranking, and include several supervised methods [140, 13, 143] mentioned in Section 5.1 for performance comparison.
The pointwise methods (Query Generation and Relevance Generation) judge the relevance of each query-document pair independently, thus offering lower time complexity and enabling batch inference. However, compared to other methods, it does not have an advantage in terms of performance. The listwise method yields significant performance especially when calling GPT-4, but suffers from expensive API cost and non-reproducibility [160]. Compared with the listwise method, the pairwise method shows competitive results based on a much smaller model FLAN-UL2 (20B). Stemming from the necessity to compare an extensive number of document pairs, its primary drawback is low efficiency.